40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 30 13:43:43 2008, rob, Update, Locking, last night 
    Reply  Wed Oct 1 00:24:54 2008, rana, Update, Locking, last night 
Message ID: 1009     Entry time: Tue Sep 30 13:43:43 2008     Reply to this: 1011
Author: rob 
Type: Update 
Category: Locking 
Subject: last night 
Steady progress again in locking again last night. Initial acquisition of DRMI+2ARMs was working well.
Short DOF handoff, CARM->MCL, AO on PO_DC, and power ramping all worked repeatedly, in the cm_step script.
This takes us to the point where the common mode servo is handed off to an RF signal and the CARM offset
is reduced to zero. This last step didn't work, but it should just require some tweaking of the gains
during the handoff.
ELOG V3.1.3-